# Short-Long Convolutions Help Hardware-Efficient Linear Attention to Focus on Long Sequences

### v1

@@ 2.1 Attention Mechanism
Given the input to the layer is $\mathbf{X}\in\mathbb{R}^{L\times d}$, where $L$ is the sequence length and $d$ is the embedding dimension, then the attention mechanism outputs pair-wise score $\mathbf{M}$
@@
Here $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ are learnable weights, and $\sigma$ is the attention function, e.g., softmax. Denote the attention matrix $\mathbf{M}\in\mathbb{R}^{L\times L}$. $\mathbf{M}_{ij}$ captures the relations between the $i$-th and the $j$-th input tokens.
@@ 2.2 Linear Recurrent Models
If we parameterize the kernel with the same length as the sequence in Eq. 5 (Fu et al., 2023b), then we can replace the SSM layer with a learned convolution kernel as a drop-in replacement. The computation can be efficiently done by using FFT theorem:
@@ 5 Experiments
To evaluate CHELA, we conduct experiments on five benchmark sequence modeling tasks across various data types, comparing them with current state-of-the-art models on each task. All experiments were realized based on NVIDIA A100-80G and Pytorch. We used float32 parameters, with bfloat16 precision for most computations.
@@ 5.1 Long-Context Sequence Modeling
We conducted experiments to evaluate sequence models using the Long Range Arena (LRA) benchmark. This benchmark, recently introduced by (Tay et al., 2020b), is designed to assess the performance of sequence models in long-context learning. The benchmark includes six tasks, namely ListOps (Nangia & Bowman, 2018), byte-level Text classification (Maas et al., 2011), byte-level document Retrieval (Radev et al., 2013), Image classification on sequences of pixels (Krizhevsky et al., 2009), Pathfinder (Linsley et al., 2018), and its extreme long version (Path-X; Tay et al. (2020b)). The input sequences for these tasks range from 1K to 16K tokens and cover various data modalities. In Table 1, CHELA is compared to various baselines, such as Transformer and its efficient versions, as well as the top-performing S4 models. To ensure a just comparison, we equalize the number of layers and model dimensions on each task so that Mega has a similar number of parameters with S4. The results are an average of five runs with distinct random seeds, and the tuning information and model details are given in the Appendix. The CHELA has shown remarkable performance over all six tasks, with an average performance of 88.26%, outperforming all the other baselines. Furthermore, we have evaluated the speed and memory efficiency of CHELA using the byte-level classification task with an input length of 4K. Our hardware-efficient linear mechanism has proven to be highly efficient, being 5.8 times faster. With its short-long convolutions hybrid design, it is worth noting that the CHELA layer is even more efficient than various linear Transformers, Structured State Space Models, and recent hybrid models by a large margin.
@@ 5.2 Raw Speech Classification
We aim to assess CHELA’s ability to model speech signals over a long range by utilizing it to classify unprocessed speech signals that have a length of 16000 instead of relying on conventional preprocessing methods, such as converting them to MFCC features. As per Gu et al. (2021) approach, we conduct speech classification on the Speech Commands dataset’s SC10 subset, which was introduced by Warden (2018). As reported in (Ma et al., 2022), the Mega-chunk uses a chunk size of 1000 to enable processing the data. As shown in Table 2, our CHELA model with 493K parameters can achieve an accuracy of 97.98%, which is the state-of-the-art method in this table. This is mainly because the large number of continuous and low-frequency signals in speech are well-suited to be processed using long convolutions. Coupled with short convolutions’ ability to rich global information, attention can easily focus on importance.
@@ 5.3 Auto-Regressive Language Modeling
By following Ma et al. (2022); Lingle (2023), we evaluate CHELA on two established language modeling benchmarks, i.e., WikiText-103 (Merity et al., 2016) and enwik8 (Mahoney, 2011), which are next-token prediction tasks. WikiText-103 is a word-level language modeling dataset containing 103M training tokens from Wikipedia articles. Following previous work (Baevski & Auli, 2019), our approach involves the utilization of adaptive softmax and input embeddings, and we work with a vocabulary of 260K tokens. Enwik8 is a widely used benchmark for character-level language modeling, which poses a significant challenge to models. It consists of around 100 million unprocessed tokens from Wikipedia articles, with a vocabulary size of about 200. To test the effectiveness of language models, we divide the test data into segments and sequentially process each segment during testing. In Table 4, we compare with previous top-performing models that are designed to take advantage of longer context, including Transformers (Baevski & Auli, 2019), Transformer-XL and S4 (Gu et al., 2021). The CHELA model has shown impressive performance on both WikiText-103 and enwik8 datasets, surpassing the baseline models by a considerable margin. Our model boasts a remarkably almost 10 times faster inference speed when compared to the Pure Transformer model. The hybrid design of the short-long convolutions layer plays a pivotal role in enabling our model to handle length extrapolation at inference time, thereby allowing it to process longer sequences than the ones seen during training. This unique feature of our CHELA enhances its ability to naturally handle complex tasks, making it a valuable asset to any long-sequence project.
@@ 5.4 Pixel-Level Sequential Image Classification
Let’s begin with tasks related to image classification, where images are treated as a sequence of pixels in one dimension. In such tasks, models are not allowed to use any two-dimensional bias from the image. Therefore, the model should be capable of identifying patterns at various timescales, including pixels that are close to the original image but far from each other in its sequence representation. We assess our model’s performance on the Sequential CIFAR-10 dataset, which is widely used as a standard benchmark for modeling long-term dependencies in RNNs. The CIFAR-10 dataset is commonly used in machine learning for image classification tasks. In this dataset, the standard train and test split is followed, where 10% of the training set is reserved for validation purposes. To classify the images, the mean of all tokens in the output sequences is calculated, and the resulting values are passed through a fully connected layer to generate class logits. The Table 3 displays the results. Our model has achieved state-of-the-art performance and the best test accuracy on the sequence classification task, surpassing numerous strong competitors such as Transformers (Vaswani et al., 2017), RNNs, state space models, and other convolutional models. Specifically, the CHELA model has outperformed the previous convolution-based models by more than ten percentage points. It is noteworthy to mention that our model has achieved impressive results by surpassing the previously established performance benchmark despite utilizing a relatively simple architecture. The model primarily uses a hybrid approach that compresses long historical information based on the output of short-long convolutions. Our most effective model comprises ten CHELA blocks, which play a significant role in achieving exceptional performance.
@@ 7 Related Works
There are various attempts to reduce the quadratic time and space complexity of vanilla attention. One approach is to employ sparse attention. That is, each token only attends to a subset of all the tokens according to pre-defined patterns, e.g., neighboring tokens within a fixed-size window. Some examples include Sparse Transformer (Child et al., 2019), BlockBERT (Qiu et al., 2020), Longformer (Beltagy et al., 2020), ETC (Ainslie et al., 2020), BigBird (Zaheer et al., 2020), HEPOS (Huang et al., 2021), and Poolingformer (Zhang et al., 2021). Another approach is to use low-rank projection (Wang et al., 2020a). Similar methods include Nyströmformer  (Xiong et al., 2021), Synthesizer (Tay et al., 2021), Transformer-LS (Zhu et al., 2021), and Luna  (Ma et al., 2021). However, these approaches face difficulty when handling causal tasks, such as auto-regressive language modeling. Kernel-based approaches can be used to approximate the full attention $\mathrm{Attn}(\mathbf{X})$. In these approaches, the quadratic-time softmax attention is replaced by fast linear-time kernel approximations (e.g., Gaussian and arc-cosine kernel). Some examples include Linear Transformer (Katharopoulos et al., 2020), Performer (Choromanski et al., 2020), Random Feature Attention (Peng et al., 2021), and FMMformer (Nguyen et al., 2021). Both low-rank projection and kernel-based approaches approximate the full attention, and thus, they often suffer from non-negligible approximation error. We can adopt clustering-based approaches, where we divide $\mathbf{Q}$ or $\mathbf{K}$ into several clusters, and perform inter-cluster attention. Such methods include Reformer (Kitaev et al., 2020), Clusterformer (Wang et al., 2020b), Sinkhorn Transformer (Tay et al., 2020a), Fast Transformer (Vyas et al., 2020), Routing Transformer (Roy et al., 2021), and FLASH (Hua et al., 2022).
@@ 8 Conclusion and Limitations
However, CHELA also has some limitations. The optimal combinations of short convolutions are not explored in this paper, which should be a future research direction of designing dynamic short-convolutional components according to the input data. Moreover, the time-varying SSM is a different idea to achieve this goal. By embracing CHELA as a starting point in the integration of hardware-efficient implementation into hybrid models, we are taking the initial step towards achieving greater efficiency gains in the future. As we continue to work towards scaling up this model to larger sizes, we are paving the way for an innovative tomorrow.

### v2

@@ 2.1 Attention Mechanism
If we have an input $\mathbf{X}$ that belongs to $\mathbb{R}^{L\times d}$, where $L$ represents sequence length and $d$ represents the embedding dimension, the attention mechanism produces pair-wise scores denoted as $\mathbf{M}$:
@@
Here $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ are learnable weights, and $\sigma$ is the attention function. Denote the attention matrix $\mathbf{M}\in\mathbb{R}^{L\times L}$. $\mathbf{M}_{ij}$ captures the relations between the $i$-th and the $j$-th input tokens.
@@ 2.2 Linear Recurrent Models
If we parameterize the kernel with the same length as the sequence in Eq. 5 (Fu et al., 2023b), then we can replace the SSM layer with a learned convolution kernel as a drop-in replacement. The computation can be efficiently done by using the FFT theorem:
@@ 5 Experiments
To assess CHELA, we performed tests on five standard sequence modeling tasks involving different types of data, and we compared them with the latest cutting-edge models for each task. All experiments were realized based on NVIDIA A100-80G and Pytorch. We used float32 parameters, with bfloat16 precision for most computations.
@@ 5.1 Long-Context Sequence Modeling
We conducted experiments to evaluate sequence models using the Long Range Arena (LRA) benchmark. This benchmark, introduced by (Tay et al., 2020b), is designed to assess the performance of sequence models in long-context learning. The benchmark includes six tasks, namely ListOps (Nangia & Bowman, 2018), byte-level Text classification (Maas et al., 2011), byte-level document Retrieval (Radev et al., 2013), Image classification on sequences of pixels (Krizhevsky et al., 2009), Pathfinder (Linsley et al., 2018), and its extreme long version (Path-X; Tay et al. (2020b)). The input sequences for these tasks range from 1K to 16K tokens and cover various data modalities. In Table 1, CHELA is compared to various baselines, such as Transformer and its efficient versions, as well as the top-performing S4 models. In order to make a fair comparison, we make sure that Mega and S4 have a similar number of parameters by balancing the number of layers and model sizes for each task. The results are based on the average of five runs with different random seeds, and you can find the tuning information and model details in the Appendix. The performance of our model has been outstanding across all six tasks, achieving an average accuracy of 88.26% and surpassing all the other comparison methods. Additionally, we assessed the speed of our model when applied to the byte-level classification task with a 4K input. Our hardware-efficient linear mechanism has demonstrated remarkable efficiency, with a speed that is 5.8 times faster. It is important to highlight that our model, with its unique short-long convolutions hybrid design, exhibits even greater efficiency compared to a variety of linear Transformers, Structured State Space Models, and recent hybrid models.
@@ 5.2 Raw Speech Classification
We intend to evaluate the capability of CHELA in modeling lengthy speech signals by employing it for the classification of unaltered speech signals with a duration of 16000, instead of depending on traditional preprocessing techniques like converting them into MFCC features. As per Gu et al. (2021) approach, we classify speech on the Speech Commands dataset’s SC10 subset, which was introduced by Warden (2018). As reported in (Ma et al., 2022), the Mega-chunk uses a chunk size of 1000 to enable processing the data. In Table 2, our model has 493K parameters and achieves a 97.98% accuracy, making it the leading method in this table. This result is primarily due to the suitability of long convolutions for processing the numerous continuous and low-frequency signals present in speech. Additionally, the ability of short convolutions to capture rich global information enables attention to focus on important aspects.
@@ 5.3 Auto-Regressive Language Modeling
By following Ma et al. (2022); Lingle (2023), we assess CHELA on two popular language modeling datasets, i.e., WikiText-103 (Merity et al., 2016) and enwik8 (Mahoney, 2011), which are next-token prediction tasks. WikiText-103 is a dataset for word-level language modeling with 103 million tokens from Wikipedia articles in its training set. In line with previous work (Baevski & Auli, 2019), our method involves using adaptive softmax and input embeddings, and we utilize a vocabulary of 260,000 tokens. Enwik8 is a commonly used benchmark for character-level language modeling, presenting a significant challenge to models. It comprises approximately 100 million unprocessed tokens from Wikipedia articles and has a vocabulary size of about 200. When evaluating language models, we segment the test data and process each segment sequentially during testing to assess their effectiveness. In Table 4, we compare with previous top-performing models that are designed to take advantage of longer context, including Transformers (Baevski & Auli, 2019), Transformer-XL and S4 (Gu et al., 2021). The model we developed demonstrated outstanding performance on both WikiText-103 and enwik8 datasets, outperforming the baseline models by a significant margin. Our model achieves an inference speed that is almost 10 times faster than the Pure Transformer model. The hybrid structure of the short-long convolutions layer plays a crucial role in enabling our model to manage length extrapolation during inference, allowing it to process longer sequences than those encountered during training. This distinctive characteristic of our model enhances its capability to naturally handle complex tasks, making it a valuable addition to any long-sequence project.
@@ 5.4 Pixel-Level Sequential Image Classification
Begin by addressing tasks related to image classification, in which images are considered as a one-dimensional sequence of pixels. In these tasks, models cannot rely on preconceived two-dimensional structures within the image. Consequently, the model must possess the ability to recognize patterns at different temporal scales, including pixels that are close together in the original image but far apart in their sequential representation. We evaluate the performance of our model using the Sequential CIFAR-10 dataset, commonly used as a benchmark for capturing long-term dependencies in RNNs. The CIFAR-10 dataset is frequently employed in machine learning for tasks on image classification. Within this dataset, the typical training and testing split is maintained, reserving 10% of the training set for validation purposes. To categorize the images, the mean of all tokens in the output sequences is computed, and the resulting values are subjected to a fully connected layer to produce class logits. The Table 3 displays the results. CHELA has achieved state-of-the-art performance and the best test accuracy on the sequence classification task, surpassing multiple strong competitors such as Transformers (Vaswani et al., 2017), RNNs, state space models, and other convolutional models. In particular, the CHELA model has exceeded the performance of previous convolution-based models by more than ten percentage points. It is important to note that our model has delivered impressive results by surpassing the previously established performance standard, even though it uses a relatively simple architecture. The model primarily employs a hybrid method that compresses long historical information based on the output of short-long convolutions. Our most effective model consists of ten CHELA blocks, which significantly contribute to achieving exceptional performance.
@@ 7 Related Works
There are various attempts to reduce the quadratic time and space complexity of vanilla attention. One approach is to employ sparse attention. That is, each token only attends to a subset of all the tokens according to pre-defined patterns, e.g., neighboring tokens within a fixed-size window. Some examples include Sparse Transformer (Child et al., 2019), BlockBERT (Qiu et al., 2020), Longformer (Beltagy et al., 2020), ETC (Ainslie et al., 2020), BigBird (Zaheer et al., 2020), HEPOS (Huang et al., 2021), and Poolingformer (Zhang et al., 2021). Another approach is to use low-rank projection (Wang et al., 2020a). Similar methods include Nyströmformer  (Xiong et al., 2021), Synthesizer (Tay et al., 2021), Transformer-LS (Zhu et al., 2021), and Luna  (Ma et al., 2021). However, these approaches face difficulty when handling causal tasks, such as auto-regressive language modeling. Kernel-based approaches can be used to approximate the full attention $\mathrm{Attn}(\mathbf{X})$. In these approaches, the quadratic-time softmax attention is replaced by fast linear-time kernel approximations (e.g., Gaussian and arc-cosine kernel). Some examples include Linear Transformer (Katharopoulos et al., 2020), Performer (Choromanski et al., 2020), Random Feature Attention (Peng et al., 2021), and FMMformer (Nguyen et al., 2021). Both low-rank projection and kernel-based approaches approximate the full attention, and thus, they often suffer from non-negligible approximation errors. We can adopt clustering-based approaches, where we divide $\mathbf{Q}$ or $\mathbf{K}$ into several clusters, and perform inter-cluster attention. Such methods include Reformer (Kitaev et al., 2020), Clusterformer (Wang et al., 2020b), Sinkhorn Transformer (Tay et al., 2020a), Fast Transformer (Vyas et al., 2020), Routing Transformer (Roy et al., 2021), and FLASH (Hua et al., 2022).
@@ 8 Conclusion and Limitations
However, CHELA also has some limitations. The optimal combinations of short convolutions are not explored in this paper, which should be a future research direction of designing dynamic short-convolutional components according to the input data. Moreover, the time-varying SSM is a different idea to achieve this goal. By embracing CHELA as a starting point in the integration of hardware-efficient implementation into hybrid models, we are taking the initial step towards achieving greater efficiency gains in the future.